ATK Tähtitieteessä harjoitustyö (syksy 2005)

Hae työohje Heikki Salolta (TÄ307; heikki.salo@oulu.fi)

ATK Tähtitieteessä harjoitustyö (syksy 2005)

Hae työohje Heikki Salolta (TÄ307; heikki.salo@oulu.fi)

ATK Tähtitieteessä harjoitustyö (syksy 2005)

Hae työohje Heikki Salolta (TÄ307; heikki.salo@oulu.fi)